A church window consisting of a rectangle topped by a semicircle is to have a perimeter p. Find the radius of the semicircle if the area of the window is to be a maximum.


1
Expert's answer
2021-12-03T07:15:16-0500

The rectangle has dimensions: "2r" and "x".

Tne semicircle: "\frac{1}{2}\cdot 2\cdot \pi \cdot r=\pi \cdot r" .

"p=2r+2x+\pi \cdot r\\
x=\frac{1}{2}(p-r(\pi+2))"

The area of the window is to be a maximum:

"S=x\cdot 2r+\frac{1}{2}\pi \cdot r^2=\\
=(p-r(\pi+2))\cdot r+\frac{1}{2}\pi \cdot r^2=\\
=pr-r^2(\pi+2)+\frac{1}{2}\pi \cdot r^2\\
S'=p-2r(\pi+2)+\pi r=p-\pi r-4r\\
p-\pi r-4r=0\\
r(\pi+4)=p\\
r=\frac{p}{\pi+4}"

If "r\in (-\infty, \frac{p}{\pi+4}), S'>0" .

If "r\in (\frac{p}{\pi+4},\infty), S'<0" .

"S(\frac{p}{\pi+4})" - max.

"r=\frac{p}{\pi+4}" .